FEL Upgrade Project Weekly Brief
July 16-20, 2001

During FEL operations this week we analyzed the terahertz radiation emanating from the recirculation bend magnets.  The quantity of terahertz radiation (~ watts) will be very useful for experiments and its signature will be a useful measure of the electron bunch characteristics.

With the cooperation of the Jefferson Lab DOE Site Office we prepared a draft interagency agreement for the AF funded UV FEL project.  We have been notified by AFRL and OSD that the funds for the first phase (FY01-02) effort for this project should be available by the end of the month.

The procurement package for the first dipole system for the Upgrade project (GW’s) were released this week.

WBS 4 (Injector):
The cathode stalk and cathode support tube drawings are in sign off. The gun chamber drawings are waiting for the alignment groups input but otherwise completed.

WBS 5 (SRF):
Receipt inspection of the first vacuum vessel and space frame continue.  Minor discrepancies have been noted including some marks on o-ring sealing surfaces and general cleanliness. Complete inspection is scheduled for completion next week.

Project engineers visited the End Can and Thermal Shield vendors.  Progress continues on schedule with no significant technical issues.

WBS 8 (Instrumentation):
User Support provided: James - one day; Rich - two days on shift.
Prototypes were completed for the Timing and Delay board and the associated "mother" board.  These will be tested next week.
A new intercom system is being installed in each of the user labs, the OCR and the Drive Laser Clean Room.  This should aid in communications and reduce noise in the control room from ringing phones.
Two new Pyro-Electric BPM detectors were wired and are awaiting testing.
Testing continues on the OBPM design.
Have started working on new Ion Pump Power Supplies for the SNS project.
Patrick and Mike have been working on the website and the Rabbit project respectfully.
Drawings completed:  OBPM/Picomotor Test Box Wiring, Assembly and Fabrication.
Check Plots Received:  Picomotor Relay Wiring Diagram, Molectron Quad OBPM Amplifier Schematic, Fabrication Assembly and Artwork and  Absolute Value/Sample & Hold Board Schematic, Assembly, Fabrication and Artwork.
Drawings at EECAD:    Analog/Differential Driver Board Schematic; Pyro-Electric Detector Buffer Amplifier Board Schematic, Assembly, Fabrication and Artwork; 32 Channel Sample and Hold Board Schematic, Assembly, Fabrication and Artwork; Ion Pump Power Supply Controller Schematic; User Lab Shutter Control Box Wiring, Schematic, Assembly and Fabrication; MPS System Drawing; Generic Temperature Monitoring Chassis Schematic, Wiring, Fabrication and Assembly.

WBS 9 (Transport):
Optical Chicane Dipoles (GW)
o  The manufacturing specifications, the drawings and the requisition were signed and sent to Procurement
    for this style of magnets.
Injector Dipoles (DU/DV)
o  The DULY continues working on the revisions to the drawings of the Injector Dipoles (GU & GV).
Arc Dipoles (GY, GX, GQ)
o  AES continued design of the Reverse Bend (GQ).  Coil leads, hoses, manifolds and cover layout and
    details were completed and the field clamps started.
o  In magnetic modeling, the final configuration of the Bend Dipole (GX) was achieved that fulfilled David
    Douglas’specifications.  The bulk field was achieved over 67% of the magnetic length to one part in
    10,000 (This worked in the IR Demo) and the current to do this will serve as the anchor current for the
    Dipole strings.  The effective length was achieved, shortening the iron length by 1 cm.
o  Additional manipulation of the magnetic model showed that moving the field clamps by 3/8 inch is the
    equivalent of a 0.5 change in iron length.
QX (3 inch quad)
o  The core contract was placed with New England Techni-Coil, a quality vendor who has supplied us in
    the past.
QT (Trim Quad)
o  The interface of the magnet to the measurement stand and the details of the advantageous use of the
    known dimension conductor that are being worked.
o  Robin Wines has not been able to continue with magnetic modeling over the last reporting period
    because of checking of the magnetic quality, specifications and drawings of the GW dipole package
o  DULY Research is working on a response to the task order form magnetic modeling and design.
Beam Line and Vacuum
o  Work continued on the Interface Control Document for the task definition for the Arc and Optical
    Chicane chambers.  The work concentrated on the specialty chambers through which the optical beam
    has to traverse.
o  Work continued on the drawing set for the girders between Cryomodules.

WBS 10 (Wiggler):
Analysis of dispersion correction coils around the entire magnet and around just the center pole indicates that the field error arises from a 3D solenoid component caused by excess reluctance on the end poles.  This is best compensated with a solenoidal coil around the whole magnet.  The dispersion section strength is unchanged when this coil is used while the coil around a single pole predominantly changes the dispersion section strength and changes the steering as a second order effect.  We have therefore settled on a design for a 3 turn correction coil around the entire magnet.  Michael Necaise is working on this.  He will also wire up the wiggler correction coils, which are now mounted on the wiggler.

The parts for the dispersion section vacuum chamber are expected in today for inspection prior to welding them up next week.  The chamber might be available for the final measurements of the dispersion section to compare the field with and without the chamber.

WBS 11 (Optics):
We are detailing the internal hardware in the optical cavity chambers, specifically: (1) edge-cooled mirror holders for the outcouplers, and the modification to the mirror mounts to accommodate them (2) y-motion of the entire mirror frame.

We completed preparations for our first use of the mirror test stand (MTS) to measure the absorption of the our 6 micron high reflectors.  Thanks to I & C for installing the RTDs and wiring the feedthrough to allow monitoring while under vacuum.

Support for operations/experimenters:
Provided optics hardware and support to M. Kelley (College of William and Mary), to E. Gillman (JLab/NSU) and D. Ermer (MSU).  Changed steering and condensing optics to mirror test stand for better clearance. Installed laser interferometer for use with MTS.  Added "tent" over optics to purge space.  Installed pyroelectric detector to calibrate total power into FIR spectrometer (for JLab THz experiment).  Checked linearity of detectors (InGaAs and Si) for autocorrelation using two-photon absorption.

Checked hardware in preparation for FEL noise (amplitude and phase) measurements to be done next week.  Changed mirrors in order to shift operation from 6 microns to 5 microns.  Provided support and optics to measure reflectivity of one of our 3 micron HR mirrors.  Inspected and passed half of our new 1.6 microns optics.

We operated the FEL all week for various users.  Work at the beginning of the week involved testing of polymer surface modification on a resonant absorption line.  We were able to run short macropulses of 4 microseconds at 20 kHz to maximize energy transfer rate.   Parasitically we were able to run X-ray measurements looking for the low energy X-rays produced while lasing around 6 microns.

We then spent two days successfully making measurements on non-linear absorption in dielectrics.  All the samples available were exposed and brought off for analysis.  We then performed alignment tests for near field IR microscopy and performed saturation tests for Beam Position Monitor electronics.   We installed a calibrated pyroelectric detector and were able to measure the total collective edge radiation power
( ~0.2 W) in the THz regime through a 1 cm diamond window. Since this is produced in short pulses and is a very intense tunable source at this wavelength it opens the possibility of a number of future scientific tests.  X-ray measurements continued through the week but struggled with signal level due to the high
background and large window absorption at the 3 keV energy range. The final day was spent measuring absorption in 6 micron mirror coatings.  Preliminary results show the losses are very small, less than 100 PPM.